# Rethinking Pseudo-Label Guided Learning for Weakly Supervised Temporal Action Localization from the Perspective of Noise Correction

### v1

@@ Methods
Influence of CALA. Experiment b incorporates the CALA algorithm, leading to a 1.3% increase in average mAP. This highlights CALA’s effectiveness in augmenting pseudo-labels through contextual information.
Influence of Customized Teacher-Student Training. In experiment c, we added the customized teacher-student framework without pseudo-label correction modules like AIC and MIC. The teacher model’s pseudo-labels guided the student model, resulting in a 0.7% performance gain. This improvement stems from the robustness of the student model, which benefits from multiple historical teacher-student iterations, reducing the impact of noisy labels.
Influence of AIC. Experiment d introduces the AIC module, yielding a 2.2% performance boost. AIC corrects inaccurately localized pseudo-labels, enhancing the learning process of the student model.
@@
Influence of MIC. Experiment e adds the MIC module, resulting in a 1.8% performance improvement. MIC compensates for missing action instances in the pseudo-labels, optimizing the student model’s learning.
Influence of Combined MIC and AIC. Experiment f applies both AIC and MIC, leading to a 4.1% improvement over experiment c. AIC rectifies overly complete instances, while MIC compensates for adjacent missing actions, effectively addressing the one-to-many problem.
Influence of High-Quality Pseudo-Label Mining. Finally, experiment g introduces high-quality pseudo-label mining loss, resulting in a 0.2% performance boost. This demonstrates its effectiveness in extracting reliable information while suppressing noise, achieving the highest performance of 42.1% mAP.
@@
It’s worth noting that our noise correction framework is decoupled from the weakly supervised models, making it applicable to any existing weakly supervised temporal action localization method. To analyze the generalizability of our noise correction framework, we apply our noise correction method sequentially to the UM (Lee et al. 2021) and CO2-Net (Hong et al. 2021), and the result is shown in Table 4.
With only the UM model, the average mAP on the THUMOS14 dataset is 30.3%. Introducing the CALA module yielded an improvement of 1.3% compared to the baseline. The addition of the Teacher-Student module contributes to the improvement of 2.0% over the baseline. When introducing the MIC, AIC, and HPM modules, the performance gains relative to the baseline are 3.6%, 5.8%, and 6.1%, respectively. Similarly, when using only the CO2-Net model, the average mAP on the THUMOS14 dataset is 36.0%. Upon sequentially introducing the CALA, Teacher-Student, MIC, AIC and HPM modules, they led to performance gains of 1.0%, 1.7%, 4.2%, 5.8%, and 6.0% compared to baseline, respectively. These results provide strong evidence for the generalizability of our noise correction framework, which can be applied to various weakly-supervised models.

### v2

@@ Methods
Influence of CALA. Experiment b applies CALA, achieving a 1.3% mAP increase, demonstrating its ability to enhance pseudo-labels with contextual information.
Influence of Customized Teacher-Student Training. Experiment c introduces a customized teacher-student framework without pseudo-label correction modules. The teacher’s labels guide the student, improving performance by 0.7% due to the student model’s robustness from prior teacher-student iterations.
Influence of AIC. Experiment d adds AIC, boosting performance by 2.2%. AIC corrects inaccurately localized pseudo-labels, refining the student model’s learning.
Influence of MIC. Experiment e incorporates MIC, resulting in a 1.8% improvement. MIC compensates for missing action instances in pseudo-labels, enhancing learning.
Influence of Combined MIC and AIC. Experiment f applies both MIC and AIC, yielding a 4.1% performance increase over experiment c. AIC corrects overly complete instances, while MIC fills in missing actions, addressing the one-to-many issue.
Influence of High-Quality Pseudo-Label Mining. Experiment g adds high-quality pseudo-label mining loss, providing a 0.2% performance boost and reaching the highest mAP of 42.1%.
@@
Our noise correction framework is decoupled from weakly supervised models, making it applicable to any existing temporal action localization method. To evaluate its generalizability, we apply our framework to the UM (Lee et al. 2021) and CO2-Net (Hong et al. 2021) models, with results shown in Table 4.
For the UM model, the average mAP on THUMOS14 is 30.3%. The CALA module improves this by 1.3%, and the Teacher-Student module contributes an additional 2.0%. The MIC, AIC, and HPM modules further increase the performance by 3.6%, 5.8%, and 6.1%, respectively.
Similarly, for CO2-Net, the mAP starts at 36.0%. The sequential introduction of the CALA, Teacher-Student, MIC, AIC, and HPM modules results in improvements of 1.0%, 1.7%, 4.2%, 5.8%, and 6.0%, respectively.
These results highlight the strong generalizability of our noise correction framework across various weakly-supervised models.